1.104 of impure sample of NaNO2 was acidified and analysed using excess iodide ion. Then I3- was titrated with 0.300 M Na2S2O3
If the titre volume is 29.25cm cube, calculate the % of NaNO2 in the sample
1
Expert's answer
2019-04-01T07:49:42-0400

Task:

1.104 of impure sample of NaNO2 was acidified and analysed using excess iodide ion. Then I3- was titrated with 0.300 M Na2S2O3. If the titre volume is 29.25 cm3, calculate the % of NaNO2 in the sample


Solution:

The amount of NaNO2 in a sample can be determined by acidifying the solution to form nitrous acid, HNO2, allowing the nitrous acid to react with an excess of iodide ion, and then titrating the triiodide ion (I3) in the resulting solution with thiosulfate (S2O32–) solution.

NaNO2 + H+ = HNO2 + Na+

The balanced redox equations are:

(1) 2H+ + 2HNO2 + 3I- = 2NO + I3- + H2O

(2) I3- + 2S2O32- = 3I- + S4O62-

According to the reaction equations:

n(HNO2) / 2 = n( I3-) = n(S2O32-) / 2

n(HNO2) = n(S2O32-) = C(S2O32-)*V(S2O32-);

n(HNO2) = C(S2O32-)*V(S2O32-) = 0.300 M * 0.02925 L = 0.008775 mol;


n(NaNO2) = n(HNO2) = 0.008775 mol;

n(NaNO2) = m(NaNO2) / M(NaNO2)

M(NaNO2) = Ar(Na) + Ar(N) + 2*Ar(O) = 23 + 14 + 2+16 = 69 g/mol;

m(NaNO2) = n(NaNO2) * M(NaNO2) = 0.008775 mol * 69 g/mol = 0.6055 g


w(NaNO2) = [m(NaNO2) / m(sample)]*100% = [0.6055 / 1.104] * 100% = 54.85%

w(NaNO2) = 54.85%


Answer: 54.85% NaNO2 in the sample.